MODE-SELECTIVITY IN THE AUTOIONIZATION DYNAMICS OF HCO

\underline{HARTMUT G. HEDDERICH}, ERIC E. MAYER and EDWARD R. GRANT, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

Rotationally resolved autoionization spectra of the HCO radical in high-Rydberg states (n = 12 − 50) built on core vibrational states, (010), (020), (030), and (001) have been obtained from rovibrational selected levels of the 3pπ^{2}Π Rydberg state. Autoionization dynamics for bending and stretching modes are analyzed and compared.